OpenStudy (anonymous):

x^3-14x=0 How many rational solutions? How many irrational solutions?

OpenStudy (anonymous):

@Directrix @jim_thompson5910 @robtobey Help please1 :)

OpenStudy (misty1212):

hey!!

OpenStudy (misty1212):

this has to factor right? since there is a common factor of \(x\) in both terms !

OpenStudy (anonymous):

Hi !

OpenStudy (anonymous):

Your right

OpenStudy (misty1212):

once you write \[x(x^2-14)=0\] it should be easy to solve (easy enuf)

OpenStudy (anonymous):

I got 3 answers

OpenStudy (misty1212):

yes, you should \[\{0,\sqrt{14},-\sqrt{14}\}\]

OpenStudy (anonymous):

There should be three though right because the leading coefficient had an exponent of 3

OpenStudy (anonymous):

Yup thats what i got

OpenStudy (anonymous):

How many are rational and how many are irrational?

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

0 is rational for sure right?

OpenStudy (misty1212):

and since \(14\) is NOT a perfect square, then \(\sqrt{14}\) is irrational

OpenStudy (anonymous):

Okay so 1 rational and 2 irrational

OpenStudy (misty1212):

yup

OpenStudy (anonymous):

Thank you so much !

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]